OpenStudy (anonymous):

can anyone help me!!!!! 6. The Martians ask you to explain one last thing, Ultimate Math Ambassador. Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will always result in the same number. You will use the inverse function that you created in problem number 5 for g(x).

OpenStudy (happinessbreaksbones):

hah I'm working on the same thing.

OpenStudy (happinessbreaksbones):

if you help me with #2 I can try to help you with this http://openstudy.com/study#/updates/5231d265e4b03eb771a2924e

OpenStudy (phi):

Do you have a g(x) function ?

OpenStudy (happinessbreaksbones):

Let g(x)=3x+1 for instance. f(g(x))=2(3x+1)+3=6x+5. g(f(x))=3(2x+3)+1=6x+10. No, they do not have to result in the same number. If we remove the m value from y=kx+m with arbitrary k values, however, they will. The same goes for f(x)=x^a and g(x)=x^b. someone asked the exact same question and got this result.. hope this helped?

OpenStudy (anonymous):

that does help thanks

OpenStudy (happinessbreaksbones):

@phi and @ganeshie8 can you guys verify if what I wrote is correct

OpenStudy (phi):

I am guessing that you have to make up a function f(x) Did you do that ? then you have to find the inverse function of f(x), which you call g(x) Did you do that ?

OpenStudy (happinessbreaksbones):

yes we had to make a function (I'm doing the same problem coincidentally)

OpenStudy (phi):

OK, as this is stephie's question, what is her f(x) and g(x) ?

OpenStudy (happinessbreaksbones):

hey I see you disappeared. If you still need help, phi went through problems #1-6 on my question. http://openstudy.com/study#/updates/5231d265e4b03eb771a2924e thanks again for your help and I hope you got the help you needed.